3. Let X be a continous random varialile. a. What valoe of e will make () valil deay? [부 ]: .. b, what is P(X 1)? mu 3% c cona..suo dus.ly fundin c. Find E(x). d. What is P(0.5 < X <1.5)?S o与 o.s =| 0,2363
Answer #1

a)

for above to be valid: \int_{0}^{2} f(x) dx =1

\int_{0}^{2} f(x) dx =\int_{0}^{2} cx4 dx = cx5/5|20 =c*32/5 =1

c=5/32 =0.15625

b) P(X=1) =0  

c)E(X) =\int_{0}^{2} xf(x) dx =\int_{0}^{2} 0.15625x5 dx = 0.15625*x6/6|20 =0.15625*64/6 =1.6667

d)

P(0.5<X<1.5)=\int_{0.5}^{1.5} f(x) dx =\int_{0.5}^{1.5} 0.15625x4 dx = 0.15625 x5/5|1.50.5 =0.2363